OpenStudy (hocuspucus292333393):

There are 20 balls in a bag. Ten of the balls are red, 4 are green, and the rest are yellow. Which ratio represents the number of yellow balls to red balls?

jimthompson5910 (jim_thompson5910):

how many are yellow?

OpenStudy (hocuspucus292333393):

is it 1:5

jimthompson5910 (jim_thompson5910):

red = 10 green = 4 yellow = y 10+4+y = 20 y = ???

OpenStudy (hocuspucus292333393):

yea its right 1/5

jimthompson5910 (jim_thompson5910):

what do you get when you isolate y in 10+4+y = 20 ?

OpenStudy (hocuspucus292333393):

@Vuriffy am I right

OpenStudy (hocuspucus292333393):

ok let me see geez

OpenStudy (hocuspucus292333393):

was I right

OpenStudy (hocuspucus292333393):

34?

jimthompson5910 (jim_thompson5910):

10+4+y = 20 14+y = 20 y = ?? you can think of it as 14+________ = 20 what must go in the blank?

OpenStudy (hocuspucus292333393):

6

jimthompson5910 (jim_thompson5910):

yes so we have 6 yellow and 10 red the ratio of yellow to red is the fraction (yellow)/(red) = 6/10 when you reduce 6/10 you get what fraction?

OpenStudy (hocuspucus292333393):

3/5

jimthompson5910 (jim_thompson5910):

yep good

OpenStudy (hocuspucus292333393):

yaya got 100% can u help with reading plz );